Determine the solution of the simultaneous nonlinear equations: x² = 5 – y² y+1 = x2

Graphically
Answer #1

Ginay x²=5-y- y to = x² 19 = x², > y = x? is a parabola upuard | and down shifted by 1 unit Do - → g*+x? $5 IS Crsicle . Contfrom Equat.com © y=x²-1 x=/+1 = + 17.5615+ = +12.5615 x = + 1.6 19 (16,1-5615) 7 0-6,1-5615) So our Solutions ao ..: (.6,1.56if you have any doubt post it